OpenStudy (anonymous):

solve by substitution help y = x + 2 -x - 2y = 8

OpenStudy (anonymous):

Cool, so what have you done so far?

OpenStudy (anonymous):

nvm I got the problem but heres another one -3x+y=-4 3x-3y=6 do I subtract 3x from -4?

OpenStudy (nincompoop):

post the question is what she's done

OpenStudy (nincompoop):

no. it says by substitution

OpenStudy (anonymous):

yeah I know but to get y =

OpenStudy (anonymous):

like y = -3x - 4?

OpenStudy (nincompoop):

substitute the first equation, that is y = x + 2 into the y of the second equation

OpenStudy (nincompoop):

y = x + 2 use this into -x - 2(x+2) = 8 this

OpenStudy (anonymous):

im talking about this problem lol -3x+y=-4 3x-3y=6

OpenStudy (anonymous):

So what you want to do is, isolate one of the variables x or y from one of the equations and substitute it into the second equation.

OpenStudy (nincompoop):

would that require substitution? maybe, but why make life difficult when you can just solve y immediately

OpenStudy (anonymous):

The questions requires to use substitution

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So which one of the equations do you want to use?

OpenStudy (anonymous):

-3x+y=-4 3x-3y=6 for this one? the top

OpenStudy (nincompoop):

-3x+y=-4 3x-3y=6 you want to use substitution for this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright, so give which variable? x or y?

OpenStudy (anonymous):

wow my english

OpenStudy (anonymous):

y

OpenStudy (anonymous):

Alright, so try isolating y on your own, and get back to me when you've done so ^.^

OpenStudy (anonymous):

okaaay well im finding x right now so 3x-3(-3x-4)=6 3x-9x+12=6? is that right so far I just need to solve it then find y

OpenStudy (anonymous):

How did you get -3x-4? Don't skip steps, just go one at a time, that way you'll learn, and eventually with enough practice you can just zoom through :P

OpenStudy (anonymous):

So, lets go back to isolating y, how would you do it?

OpenStudy (anonymous):

because -3x+y=-4

OpenStudy (anonymous):

y = -4 - 3x? Do you see the problem?

OpenStudy (anonymous):

I don't remember :( its been awhile

OpenStudy (anonymous):

 Just take it step by step with me :) and don't rush!